OpenStudy (anonymous):

What is the exact value of cos 5pi/4 radians?

OpenStudy (anonymous):

I was thinking that i would draw a triangle but i didn't know which one?

OpenStudy (raden):

5pi/4 radians = 225 degrees it very familiar in trigono

OpenStudy (anonymous):

what about the cos?

hero (hero):

pi = 180 \[\cos\left(\frac{5\pi}{4}\right) = \cos\left(\frac{5(180)}{4}\right) = \cos(5(45)) = \cos(225) = -\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

cos225 is 1/sqrt(2)

hero (hero):

@MCJones9879, cos(225) is in the 3rd quadrant so it has to be negative.

OpenStudy (anonymous):

-1/sqrt(2) i meant .. sorry

OpenStudy (anonymous):

Do i draw the triangle perpendicular to the y or x axis?

OpenStudy (anonymous):

or -sqrt(2)/2

hero (hero):

I'm not a big fan of drawing

OpenStudy (anonymous):

Does it ask you to draw a triangle or just find the answer?

OpenStudy (anonymous):

How'd you know it was - rad2/2?

OpenStudy (anonymous):

@MCJones9879 No it doesn't but I'm just wondering how he got it

hero (hero):

Unit circle bro

OpenStudy (anonymous):

Depending on what quadrant you are in it will be positive or negative

OpenStudy (anonymous):

that i now but what about the numbers?

OpenStudy (anonymous):

Since you are asking for specfically cos it is negative since like Hero said, it is in the 3rd quadrant

OpenStudy (anonymous):

you look up cos 225- on the unit circle being in the third quadrant

OpenStudy (anonymous):

cos(225)=cos(45+180) =cos45cos180+sin45sin180 =sqrt(2)/2*(-1)+sqrt(2)*(0) =-sqrt(2)/2+0 =-sqrt(2)/2

OpenStudy (anonymous):

Here is a link for quick answers for sin, cos, tan , cot, sec, and csc http://mathmistakes.info/facts/TrigFacts/learn/vals/sum.html

hero (hero):

\[\sin(0, 30, 45, 60, 90) = \frac{\sqrt{0,1,2,3,4}}{2}\] \[\cos(0, 30, 45, 60, 90) = \frac{\sqrt{4, 3, 2, 1, 0}}{2}\]

hero (hero):

If you can understand it

hero (hero):

Nobody wants to memorize a gigantic circle

OpenStudy (anonymous):

@Hero I understand it

hero (hero):

Well, that's great. The unit circle is okay to learn but there are much easier ways

OpenStudy (anonymous):

Well, 5pi/4=225, so what you have me above wouldn't apply, right? Those numbers are way too low

OpenStudy (anonymous):

or maybe I don't understand it?

hero (hero):

You have to correspond them. 225 is a multiple of 45

hero (hero):

every multiple of 45 is going to be \[±\frac{\sqrt{2}}{2}\]

hero (hero):

Compare what I wrote with the unit circle. You'll start to see some correspondences.

hero (hero):

All you need to do is memorize the first quadrant and you practically have the whole thing memorized. The only thing to pay attention to afterwards are the signs.

OpenStudy (anonymous):

yea the quadrants, but thats simple, thnx a lot for helping me, helped a lot!